5 points Save Answer s a potenial diference between the plates of 8 V lf the charge on one of the plates of the capacitor is恨。iC what is the ele tica e e y tor by this capacitor? 6.0× 10-8 J 320 x 10-6 J 50-10-8 J O 30 108 J 640-10-6 J

A thin, circular disk of radius R·30 cris oriented in the yz plane with its center as the origin. The disk carries a total charge Q Calculate the magnitude of the electric Sold due to the disk at the pointx 15 cm along the x-axis 3 p distributed uniform y over its s ace O 458 105 NC 199 105 NNC 9.95 105 NIC 249 105 NC 3 32 105 N/C
Answer #1

Q1:

electrical energy stored in the capacitor, E = (1/2) * Q*V = (0.5) * (8 *10-6) * (80) = 320 * 10-6 J

Option (ii) is the correct choice.